A lift cage of mass 500 kg accelerates upwards from rest to a velocity of 6 ms−1 whilst travelling a distance of 12 m. The frictional resistance to motion is 200 N. Making use of the principle of conservation of energy, determine:

i) the work done

ii) the tension in the lifting cable

iii) the maximum power developed.


1
Expert's answer
2021-02-24T07:05:45-0500

m = 500 kg − mass of the cage;

𝑣 = 6 m/s − final velocity of the cage;

h = 12m − distance traveled by the cage;

F = 200N − frictional resistance to motion;

t − time of travelling;

a − acceleration of the cage.


i) The lift, after accelerating h upwards gains:

- its potential energy "E_p =mgh,"

- its kinetic energy "E_k=mv^2\/2,"

- in addition its motor needs to work against friction, "W_F=Fh."

The sum of these 3 energies is equal to the driving motor's work


"W=E_p+E_k+W_F"

"=mgh+mv^2\/2+Fh"

"W=500kg(9.81m\/s^2)12m+500kg(6m\/s)^2\/2"

"+200N(12m)=70260\ J=70.26\ kJ"

ii) The whole amount of work "(W)" is done by the tension force "\vec N" (pointing upwards) on the distance "h." Because "W=N\cdot h," we've got:


"N=\dfrac{70260\ J}{12m}=5855\ N"

iii) Rate equation for the cage:


"0=\text{v}-at"

Equation of motion for the cage:


"s=\dfrac{at^2}{2}"

"s=\dfrac{(\dfrac{\text{v}}{t})t^2}{2}"

"t=\dfrac{2s}{\text{v}}"

The motor's power P equals " work / time". Therefore:


"P=\dfrac{W}{t}=\dfrac{W\text{v}}{2s}"

"P=\dfrac{70260\ J(6m/s)}{2(12m)}=17565\ W=17.565\ kW"